OpenStudy (anonymous):

I need help finding the range, y- intercept and the asymptotes for the following problem:f(x)=3^x-2

OpenStudy (cwrw238):

y intercept is when x = 0 = 3^0 - 2 = 1-2 = -1 oh - thats only correct if the exponent = x (or is it x-2? )

OpenStudy (anonymous):

I meant the x- intercept

OpenStudy (anonymous):

I got the y-intercept

OpenStudy (anonymous):

(0,-1)